The width of a rectangle is 9ft more than 3 times the length. Find the dimensions of the rectangle of the perimeter is 50ft
Found 2 solutions by addingup, ikleyn:
Answer by addingup(3677)

W = 3L+9
2L+2(3L+9) = 50
2L+6L+18 = 50
8L = 32
L = 4
W = 3(4)+9 = 21

Answer by ikleyn(52778)

.
Commonly used and commonly accepted way of using the terms "width" and "length" is this:

In a rectangle, the length is the greater of two dimensions, while the width is the smaller of two dimensions.

The given condition violates this commonly accepted way.

Therefore, in my view, the right place of this problem is in TRASH section with the WARNING issued to the author.